OpenStudy (anonymous):

What is the GCF of the numerator and denominator of the expression 4x+20/x^2+2x-15?

OpenStudy (anonymous):

Denominator is x^2+2x-15 =x^2+5x-3x-15 =x(x+5)-3(x+5) =(x+5)(x-3) Numerator is 4x+20 =4(x+5) The common factor is x+5